# Is uniform expressivity too restrictive? Towards efficient expressivity of GNNs

### v1

@@ Abstract
Uniform expressivity guarantees that a Graph Neural Network (GNN) can express a query without the parameters depending on the size of the input graphs. This property is desirable in applications in order to have number of trainable parameters that is independent of the size of the input graphs. Uniform expressivity of the two variable guarded fragment (GC2) of first order logic is a well-celebrated result for Rectified Linear Unit (ReLU) GNNs Barceló et al. (2020). In this article, we prove that uniform expressivity of GC2 queries is not possible for GNNs with a wide class of Pfaffian activation functions (including the sigmoid and $\tanh$), answering a question formulated by Grohe (2021). We also show that despite these limitations, many of those GNNs can still efficiently express GC2 queries in a way that the number of parameters remains logarithmic on the maximal degree of the input graphs. Furthermore, we demonstrate that a log-log dependency on the degree is achievable for a certain choice of activation function. This shows that uniform expressivity can be successfully relaxed by covering large graphs appearing in practical applications. Our experiments illustrates that our theoretical estimates hold in practice.
@@ 1 Introduction
Graph Neural Networks (GNNs) form a powerful computational framework for machine learning on graphs, and have proven to be very performant methods for various applications ranging from analysis of social networks, structure and functionality of molecules in chemistry and biological applications Duvenaud et al. (2015); Zitnik et al. (2018); Stokes et al. (2020); Khalife et al. (2021), computer vision Defferrard et al. (2016), simulations of physical systems Battaglia et al. (2016); Sanchez-Gonzalez et al. (2020), and techniques to enhance optimization algorithms Khalil et al. (2017); Cappart et al. (2021) to name a few. Significant progress has been made in recent years to understand their computational capabilities, in particular regarding functions one can compute or express via GNNs. This question is of fundamental importance as it precedes any learning related consideration, by asking a description the class of functions one can hope to learn with a GNN. A better understanding of the dependency on the architecture of the GNN (activation and aggregation functions, number of layers, inner dimensions, …) can also help to select the most appropriate one, given some basic knowledge on the structure of the problem at hand. One common approach in this line of research is to compare standard algorithms on graphs to GNNs. For example, if a given algorithm is able to distinguish two graph structures, is there a GNN able to distinguish them (and conversely)? The canonical algorithm that serve as a basis of comparison is the color refinement algorithm (closely related to Weisfeler-Lehman algorithm), a heuristic that almost solve graph isomorphism Cai et al. (1992). It is for example well-known that the color refinement algorithm refine the standard GNNs, and that the activation function has an impact on the the ability of a GNN to refine color refinement Morris et al. (2019); Grohe (2021); Khalife & Basu (2023).
@@
Evidently, the logic of GNNs depends on the choice of architecture. Recent work include the impact of the activation function aggregation function on uniform expressivity Khalife (2023); Rosenbluth et al. (2023). In Grohe & Rosenbluth (2024), the authors also study the expressivity depending on the type of messages that is propagated (message that depends only on the source vs vs. dependency on source and target vertex). They show that in a non-uniform setting, the two types of GNNs have the same expressivity, but the second variant is more expressive uniformly. Additionally, Khalife (2023) also showed that GNNs with rational activations and aggregations do not allow to express such queries uniformly, even on trees of depth two. More generally, a complete characterization of the activation function that enables GC2 expressivity remains elusive.
@@
By allowing the size of the GNN to grow with the input graphs, it is possible to relax the uniform notion of expressivity and ask analogous questions to the uniform case. For instance, what is the portion of first-order logic one can express over graphs of bounded order with a GNN? What is the required size of a GNN with a given activation to do so? In this setup, it becomes now possible that general GNNs express a broader class than the uniform one (GFO+C), and that $\Sigma$-AC-GNNs express a broader class than GC2. Such question has been explored in the work of Grohe (2023), that describes the new corresponding fragments of FO+C. In particular, the author proves an equivalence between a superclass of GFO+C, called GFO+$\text{C}_{\text{nu}}$ and general GNNs with rational piecewise linear (rpl)-approximable functions. More precisely, a query can be expressed by a family of rpl-approximable GNNs whose number of parameters grow polynomially with respect to the input graphs, if and only if the query belongs to GFO+$\text{C}_{\text{nu}}$, a more expressive fragment than GFO+C. The GNNs used to prove that a GFC+$\text{C}_{\text{nu}}$ this equivalence only requires linearized sigmoid activations ($x\mapsto\min(1,\max(0,x))$) and $\Sigma$-aggregation. Similarly to the uniform case, the set of activation functions that allows maximal expressivity still needs to be better understood. If one restricts to $\Sigma$-AC-GNNs and the well-known GC2 fragment, one can ask for example, given a family of activation functions, how well $\Sigma$-AC-GNNs with those activation functions can express GC2 queries. For example, given an activation function, and integer $n$, what is the number of parameters of a $\Sigma$-AC-GNNs needed to express GC2 queries of depth [^1] The notion of depth will be defined more precisely in Section 2. For the time being, one can think of the depth as a measure of complexity measure on the space of queries. $d$ over graphs of order at most $n$?
@@
Main contributions. Our first contribution goes towards a more complete understanding of both uniform and non-uniform expressivity of GNNs, with a focus on $\Sigma$-AC-GNNs. We first show that similarly to rational GNNs, there exist a large family of GC2 that cannot be expressed with $\Sigma$-AC-GNNs with bounded Pfaffian activations including the sigmoidal activation $x\mapsto\frac{1}{1+\exp(-x)}$, function that was so far conjectured to be as expressive as the ReLU $x\mapsto\max(0,x)$ units in this setup Grohe (2021). More precisely, we prove that GNNs within that class cannot express all GC2 queries uniformly, in contrast with the same ones replaced with ReLU activations. Our second contribution adds on descriptive complexity results Grohe (2023) by providing new complexity upper-bounds in the non-uniform case for a class of functions containing the Pffafian ones. Our constructive approach takes advantage of the composition power providing some simple assumptions on the activation functions, endowing deep neural networks to approximate the step function $x\mapsto\mathbf{1}_{\{x>0\}}$ efficiently. Our experiments confirm our theoretical statements on synthetic datasets.
@@ 2 Preliminaries
Notations. In the following, for any positive integer $n$, $[n]$ refers to the set $\{1,\cdots,n\}$. We assume the input graphs of GNNs to be finite, undirected, simple, and vertex-labeled: a graph is a tuple $G=(V(G),E(G),P_{1}(G),\allowbreak\cdots,P_{\ell}(G))$ consisting of a finite vertex set $V(G)$, a binary edge relation $E(G)\subset V(G)^{2}$ that is symmetric and irreflexive, and unary relations $P_{1}(G),\cdots,P_{\ell}(G)\subset V(G)$ representing $\ell>0$ vertex labels. In the following, we suppose that the $P_{i}(G)$’s form a partition of the set of vertices of $G$, i.e. each vertex has a unique label. Also, the number $\ell$ of labels, which we will also call colors, is supposed to be fixed and does not grow with the size of the input graphs. In order to describe the logic of GNNs, we also take into account access to the color of the vertices into the definition of the logic considered. When there is no ambiguity about which graph $G$ is being considered, $N(v)$ refers to the set of neighbors of $v$ in $G$ not including $v$. $\lvert G\rvert$ will denote the number of vertices of $G$. We use simple curly brackets for a set $X=\{x\in X\}$ and double curly brackets for a multiset $Y=\{\{y\in Y\}\}$.
@@
Fix an activation function $\sigma:\mathbb{R}\to\mathbb{R}$. For any number of hidden layers $k\in\mathbb{N}$, input and output dimensions $w_{0}$, $w_{k+1}\in\mathbb{N}$, a $\mathbb{R}^{w_{0}}\rightarrow\mathbb{R}^{w_{k+1}}$ neural network with $\sigma$ activation is given by specifying a sequence of $k$ natural numbers $w_{1},w_{2},\cdots,w_{k}$ representing widths of the hidden layers and a set of $k+1$ affine transformations $T_{i}:\mathbb{R}^{w_{i-1}}\rightarrow\mathbb{R}^{w_{i}}$, $i=1,\ldots,k+1$. Such a NN is called a $(k+1)$-layer NN, and is said to have $k$ hidden layers. The function $f:\mathbb{R}^{w_{0}}\rightarrow\mathbb{R}^{w_{k+1}}$ computed or represented by this NN is:
$f=T_{k+1}\circ\sigma\circ T_{k}\circ\cdots T_{2}\circ\sigma\circ T_{1}.$
The size of the neural network is defined as $w_{1}+\cdots+w_{L}$.
@@
A GNN [^2] This definition coincides with aggregation-combine GNNs (AC-GNNs) in the litterature. In the following, for simplicity, unless state otherwise, we refer to AC-GNNs simply as GNNs. is a recursive embedding of vertices of a labeled graph by relying on the underlying adjacency information and node features using a neural network (cf. Definition 2.1). Each vertex $v$ is attributed an indicator vector $\xi^{0}(v)$ of size $\ell$, encoding the color of the node $v$: the colors being indexed by the palette $\{1,\cdots,\ell\}$, $\xi^{0}(v)=e_{i}$ (the $i$-th canonical vector) if the color of the vertex $v$ is $i$. A GNN is characterized by:
$\circ$ A positive integer $T$ called the number of iterations, positive integers $(d_{t})_{t\in\{1,\cdots,T\}}$ and $(d^{\prime}_{t})_{t\in\{0,\cdots,\linebreak T\}}$ for inner dimensions. $d_{0}=d^{\prime}_{0}=\ell$ is the input dimension of the GNN (number of colors) and $d_{T}$ is the output dimension.
$\circ$ A sequence of combination and aggregation functions $(\mathsf{comb}_{t},\allowbreak\mathsf{agg}_{t})_{t\in\{1,\cdots,T\}}$. Each aggregation function $\mathsf{agg}_{t}$ maps each finite multiset of vectors of $\mathbb{R}^{d_{t-1}}$ to a vector in $\mathbb{R}^{d^{\prime}_{t}}$. For any $t\in\{1,\cdots,T\}$, each combination function $\mathsf{comb}_{t}:\mathbb{R}^{d_{t-1}+d^{\prime}_{t}}\longrightarrow\mathbb{R}^{d_{t}}$ is a neural network with given activation function $\sigma:\mathbb{R}\longrightarrow\mathbb{R}$.
The update rule of the GNN at iteration $t\in\{0,\cdots,T-1\}$ for any labeled graph $G$ and vertex $v\in V(G)$, is given by:
$\xi^{t+1}(v)=\mathsf{comb}(\xi^{t}(v),\mathsf{agg}\{\{\xi^{t}(w))\,:\,w\in\mathcal{N}_{G}(v)\}\})$
The size of a GNN is the maximum size of the combine neural network functions.
@@
  • ReLU: The Rectified Linear Unit (ReLU) activation function $\operatorname{ReLU}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is defined as $\operatorname{ReLU}(x)=\max\{0,x\}$.
  • CReLU: The Clipped Rectified Linear Unit (CReLU) activation function $\operatorname{CReLU}:\mathbb{R}\rightarrow[0,1]$ is defined as $\operatorname{CReLU}(x)=\min\{\max\{0,x\},1\}$.
  • tanh: The Hyperbolic Tangent activation function $\tanh:\mathbb{R}\rightarrow(-1,1)$ is defined as $\tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}.$
  • Sigmoid: The Sigmoid activation function $\operatorname{Sigmoid}:\mathbb{R}\rightarrow(0,1)$ is defined as $\operatorname{Sigmoid}(x)=\frac{1}{1+e^{-x}}.$
Definition 2.3 (Guarded model logic (GC) and GC2 (Barceló et al., 2020; Grohe, 2021)).
The fragment of guarded logic GC is formed using quantifiers that restrict to range over the neighbours of the current nodes. GC-formulas are formed from the atomic formulas by the Boolean connectives and quantification restricted to formulas of the form $\exists^{\geq p}y(E(x,y)\land\psi))$, where x and y are distinct variables and $x$ appears in $\psi$ as a free variable. Note that every formula of GC has at least one free variable. We refer to the 2-variable fragment of GC as GC2, also known as graded modal logic (with two variables).
A more constructive definition can be given as follows: a graded modal logic formula is either $\mathsf{Col}(x)$ (returning 1 or 0 for one of the palette colors) or one of the following:
$\lnot\phi(x),\quad\phi(x)\land\psi(x),or\quad\exists^{\geq N}y(E(x,y)\land\phi(y))$
where $\phi$ and $\psi$ are GC2 formulas and $N$ is a positive integer:
@@
All graded modal logic formulas are unary by definition, so all of them define unary queries. Suppose $\ell=2$ (number of colors), and for illustration purposes $\mathsf{Col}_{1}=\text{Red}$, $\mathsf{Col}_{2}=\text{Blue}$. Let:
$\gamma(x):=\text{Blue}(x)\land\exists y(E(x,y)\land\exists^{\geq 2}x(Edge(y,x)\land\text{Red}(x))$
$\gamma$ queries if $x$ has blue color and as at least one neigbhor which has two blue neighbors. Then $\gamma$ is in GC2. Now,
$\delta(x):=\text{Blue}(x)\land\exists y(\lnot E(x,y)\land\exists^{\geq 2}xE(y,x)\land\text{Red}(x))$
@@
$\eta(x):=\lnot(\exists y(E(x,y)\land\exists^{\geq 2}xE(y,x)\land\text{Blue}(x))$
@@
(Khovanskiĭ, 1991) Let $U\subseteq\mathbb{R}^{n}$ be an open set. A Pfaffian chain of order $r\geq 0$ and degree $\alpha\geq 1$ in U is a sequence of analytic real functions $f_{1},\cdots,f_{r}$ in U satisfying differential equations:
$\frac{\partial f_{i}}{\partial x_{j}}=P_{i,j}(x,f_{1}(x),\ldots,f_{i}(x)))$
where $P_{i,j}\in\mathbb{R}[x_{1},...,x_{n},y_{1},...,y_{i}]$ are polynomials of $(n+i)$ variables of degree at most $\alpha$.

### v2

@@ Abstract
Uniform expressivity guarantees that a Graph Neural Network (GNN) can express a query without the parameters depending on the size of the input graphs. This property is desirable in applications in order to have a number of trainable parameters that is independent of the size of the input graphs. Uniform expressivity of the two variable guarded fragment (GC2) of first order logic is a well-celebrated result for Rectified Linear Unit (ReLU) GNNs Barceló et al. (2020). In this article, we prove that uniform expressivity of GC2 queries is not possible for GNNs with a wide class of Pfaffian activation functions (including the sigmoid and $\tanh$), answering a question formulated by Grohe (2021). We also show that despite these limitations, many of those GNNs can still efficiently express GC2 queries in a way that the number of parameters remains logarithmic on the maximal degree of the input graphs. Furthermore, we demonstrate that a log-log dependency on the degree is achievable for a certain choice of activation function. This shows that uniform expressivity can be successfully relaxed by covering large graphs appearing in practical applications. Our experiments illustrates that our theoretical estimates hold in practice.
@@ 1 Introduction
Graph Neural Networks (GNNs) form a powerful computational framework for machine learning on graphs, and have proven to be very performant methods for various applications ranging from analysis of social networks, structure and functionality of molecules in chemistry and biological applications Duvenaud et al. (2015); Zitnik et al. (2018); Stokes et al. (2020); Khalife et al. (2021), computer vision Defferrard et al. (2016), simulations of physical systems Battaglia et al. (2016); Sanchez-Gonzalez et al. (2020), and techniques to enhance optimization algorithms Khalil et al. (2017); Cappart et al. (2021) to name a few. Significant progress has been made in recent years to understand their computational capabilities, in particular regarding functions one can compute or express via GNNs. This question is of fundamental importance as it precedes any learning aspect, by asking a description the class of functions one can hope to learn via a GNN. A better understanding of the dependency of the expressivity on the architecture of the GNN (activation and aggregation functions, number of layers, inner dimensions, …) can also help to select the most appropriate one, given some basic knowledge on the structure of the problem at hand. One common approach in this line of research is to compare standard algorithms on graphs to GNNs. For example, if a given algorithm is able to distinguish two graph structures, is there a GNN able to distinguish them (and conversely)? The canonical algorithm that serve as a basis of comparison is the color refinement algorithm (closely related to Weisfeler-Lehman algorithm), a heuristic that almost solve graph isomorphism Cai et al. (1992). It is for example well-known that the color refinement algorithm refine the standard GNNs, and that the activation function has an impact on the the ability of a GNN to refine color refinement Morris et al. (2019); Grohe (2021); Khalife & Basu (2023).
@@
Evidently, the logic of GNNs depends on the choice of architecture. Recent work include the impact of the activation function aggregation function on uniform expressivity Khalife (2023); Rosenbluth et al. (2023). In Grohe & Rosenbluth (2024), the authors also study the expressivity depending on the type of messages that is propagated (message that depends only on the source vs vs. dependency on source and target vertex). They show that in a non-uniform setting, the two types of GNNs have the same expressivity, but the second variant is more expressive uniformly. Additionally, it was also showed that GNNs with rational activations and aggregations do not allow to express such queries uniformly, even on trees of depth two Khalife (2023). More generally, a complete characterization of the activation function that enables GC2 remains elusive.
@@
By allowing the size of the GNN to grow with the input graphs, it is possible to relax the uniform notion of expressivity and ask analogous questions to the uniform case. For instance, what is the portion of first-order logic one can express over graphs of bounded order with a GNN? What is the required size of a GNN with a given activation to do so? In this setup, it becomes now possible that general GNNs express a broader class than the uniform one (GFO+C), and that $\Sigma$-AC-GNNs express a broader class than GC2. Such question has been explored in the work of Grohe (2023), that describes the new corresponding fragments of FO+C. In particular, the author proves an equivalence between a superclass of GFO+C, called GFO+$\text{C}_{\text{nu}}$ and general GNNs with rational piecewise linear (rpl)-approximable functions. More precisely, a query can be expressed by a family of rpl-approximable GNNs whose number of parameters grow polynomially with respect to the input graphs, if and only if the query belongs to GFO+$\text{C}_{\text{nu}}$, a more expressive fragment than GFO+C. The GNNs used to prove this equivalence only requires linearized sigmoid activations ($x\mapsto\min(1,\max(0,x))$) and $\Sigma$-aggregation. Similarly to the uniform case, the set of activation functions that allows maximal expressivity still needs to be better understood. If one restricts to $\Sigma$-AC-GNNs and the well-known GC2 fragment, one can ask for example, given a family of activation functions, how well $\Sigma$-AC-GNNs with those activation functions can express GC2 queries. For example, given an activation function, and integer $n$, what is the number of parameters of a $\Sigma$-AC-GNNs needed to express GC2 queries of depth [^1] The notion of depth will be defined more precisely in Section 2. For the time being, one can think of the depth as a measure of complexity measure on the space of queries. $d$ over graphs of order at most $n$?
@@
Main contributions. Our first contribution goes towards a more complete understanding of both uniform and non-uniform expressivity of GNNs, with a focus on $\Sigma$-AC-GNNs. We first show that similarly to rational GNNs, there exist a large family of GC2 that cannot be expressed with $\Sigma$-AC-GNNs with bounded Pfaffian activations including the sigmoidal activation $x\mapsto\frac{1}{1+\exp(-x)}$, function that was so far conjectured to be as expressive as the ReLU $x\mapsto\max(0,x)$ units in this setup Grohe (2021). More precisely, we prove that GNNs within that class cannot express all GC2 queries uniformly, in contrast with the same ones replaced with ReLU activations. Our second contribution adds on descriptive complexity results Grohe (2023) by providing new complexity upper-bounds in the non-uniform case for a class of functions containing the Pffafian ones. Our constructive approach takes advantage of the composition power, provided some simple assumptions on the activation functions, endowing deep neural networks to approximate the step function $x\mapsto\mathbf{1}_{\{x>0\}}$ efficiently. Our experiments confirm our theoretical statements on synthetic datasets.
@@ 2 Preliminaries
Notations. In the following, for any positive integer $n$, $[n]$ refers to the set $\{1,\cdots,n\}$. We assume the input graphs of GNNs to be finite, undirected, simple, and vertex-colored with $\ell\geq 1$ colors: a graph is a tuple $G=(V(G),E,\lambda)$ consisting of a finite vertex set $V(G)$, a binary edge relation $E\subset V(G)^{2}$ that is symmetric and irreflexive, representing the edges; and a map $\lambda:V(G)\rightarrow[\ell]$, representing the colors (or labels). The number of colors $\ell$ is fixed and does not depend on $n$. In $G$, $\mathcal{N}_{G}(v)$ will denote the set of vertices adjacent to $v$ in $G$ (not including $v$).
Given a function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$, $\sigma^{\prime}$ will denote the derivative and $\sigma^{N}=\sigma\circ\cdots\circ\sigma$ (i.e., $\sigma^{1}=\sigma$ and $\sigma^{N}=\sigma^{N-1}\circ\sigma$) the $N$-th iterated composition of $\sigma$. The function $\log$ refers to the logarithm in base 2.
@@
A $(k+1)$-layer neural network (NN) with activation function $\sigma:\mathbb{R}\to\mathbb{R}$ with input dimension $w_{0}$ and output dimension $w_{k+1}$ is formed by a sequence of $k+1$ affine transformations
$T_{i}:\mathbb{R}^{w_{i}}\rightarrow\mathbb{R}^{w_{i+1}}\qquad(i\in\{0,\ldots,k\})$
where $w_{0},\ldots,w_{k+1}$ are positive integers. The function that the NN computes is the function $f:\mathbb{R}^{w_{0}}\rightarrow\mathbb{R}^{w_{k+1}}$ given by
$f=T_{k}\circ\sigma\circ T_{k-1}\circ\cdots T_{1}\circ\sigma\circ T_{0}$
where $\sigma$ is applied pointwise for every coordinate of each inner layer’s output. The size of the neural network is $w_{0}+w_{1}+\cdots+w_{k}+w_{k+1}$.
@@
A GNN [^2] This definition coincides with sum-aggregation-combine GNNs ($\Sigma$-AC-GNNs) in the literature. In the following, for simplicity, unless state otherwise, we refer to $\Sigma$-AC-GNNs simply as GNNs. Some of our results generalize to a slightly more general class of aggregation functions, see Remark B.1 in the appendix. is a recursive embedding of vertices of a graph. At each iteration, the GNN combines the information at the current vertex as well as the aggregated information on the vertex’s neighborhood using a neural network. More precisely, a GNN with $T$ iterations, activation function $\sigma:\mathbb{R}\longrightarrow\mathbb{R}$ and input size $\ell$ is a sequence of $T$ combination functions
$\mathsf{comb}_{t}:\mathbb{R}^{d_{t}}\times\mathbb{R}^{d_{t}}\rightarrow\mathbb{R}^{d_{t+1}}\qquad(t\in\{0,\ldots,T-1\})$
that are NNs with activation function $\sigma$, and where $d_{0}=\ell$. Given a graph $G$ with colors $1,\ldots,\ell$, the GNN builds $T+1$ vertex embeddings $\xi^{t}(v,G)\in\mathbb{R}^{d_{t}}$ ($t\in\{0,\ldots,T-1\}$, $v\in V(G)$) as follows:
$\circ$ The initial embedding $\xi^{0}(v,G)$ is given by $\xi^{0}(v,G)=e_{\lambda(v)}$ where $e_{i}\in\mathbb{R}^{\ell}$ is the $i$th canonical vector and $\lambda(v)\in[\ell]$ the label of $v$.
$\circ$ At iteration $t$, $\xi^{t+1}(v,G)$ is computed from $\xi^{t}(v,G)$ via the update rule
$\xi^{t+1}(v,G)=\mathsf{comb}_{t}\left(\xi^{t}(v,G),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{t}(w,G)\right).$ (1)
The size of the GNN is the maximum size of the NNs given by the combination functions.
@@
  • ReLU: The Rectified Linear Unit $\operatorname{ReLU}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is defined as $\operatorname{ReLU}(x)=\max\{0,x\}$.
  • CReLU: The Clipped Rectified Linear Unit $\operatorname{CReLU}:\mathbb{R}\rightarrow[0,1]$ defined by $\operatorname{CReLU}(x)=\min\{\max\{0,x\},1\}$.
  • tanh: The Hyperbolic Tangent $\tanh:\mathbb{R}\rightarrow(-1,1)$ given by $\tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}.$
  • Sigmoid: The Sigmoid $\operatorname{Sigmoid}:\mathbb{R}\rightarrow(0,1)$ defined by $\operatorname{Sigmoid}(x)=\frac{1}{1+e^{-x}}.$
Definition 2.3 (GC2 (Barceló et al., 2020; Grohe, 2021)).
A GC2 query is a formula with one free variable obtained from the always true empty formula t and atomic formulas $\mathsf{Col}(x)$ (returning 1 or 0 for one of the palette colors) through one of the following operations
$\lnot\phi(x),~{}~{}\phi(x)\land\psi(x)~{}~{}\text{ and }~{}~{}\exists^{\geq N}y(E(x,y)\land\phi(y))$
where $\lnot$ is the logical negation, $\land$ the logical conjunction, $\exists^{\geq N}$, with $N$ a positive integer, means “there exist at least $N$”, and $E(x,y)$ means “$x$ and $y$ are adjacent in the considered graph”. The depth, $\mathop{\mathrm{D}}(\,\cdot\,)$ is a complexity measure of GC2 queries defined recursively as follows: for the always true empty formula t, $\mathop{\mathrm{D}}(\textsc{t})=0$; for any atomic formula, $\mathop{\mathrm{D}}(\mathsf{Col}(\cdot))=1$; and for the non-atomic formulas, $\mathop{\mathrm{D}}(\lnot\phi)=\mathop{\mathrm{D}}(\phi)+1$, $\mathop{\mathrm{D}}(\phi\land\psi)=\mathop{\mathrm{D}}(\phi)+\mathop{\mathrm{D}}(\psi)+1$ and $\mathop{\mathrm{D}}\!\left(\exists^{\geq N}y(E(x,y)\land\phi(y))\right)=\mathop{\mathrm{D}}(\phi)+1$.
Remark 2.1.
GC2 queries are formulas from the so-called guarded model logic (GC), but restricted with two variables. In general, GC formulas are free formulas constructed using Boolean connectives, such as $\lnot$ and $\land$, and quantifiers that range over the neighbors of the current node, such as $\exists^{\geq N}y(E(x,y)\land\phi(y))$.
Remark 2.2.
Given our definition, the depth of a query depends on its writing. For example, if $\phi$ is a GC2 query, $\phi$ and $\phi\land\phi$ express the same GC2 query, but the latter has larger depth.
Definition 2.4.
Given a graph $G$, a vertex $v$ and a unary query $Q$, $Q(v,G)=1$ if $Q(v)$ is true in $G$ and $Q(v,G)=0$ if $Q(v)$ is false in $G$.
@@
All GC2 formulas define unary queries. Suppose $\ell=2$ (number of colors), and for illustration purposes $\mathsf{Col}_{1}=\text{Red}$, $\mathsf{Col}_{2}=\text{Blue}$. Then
$\gamma(x):=\text{Blue}(x)\land\exists y\left(E(x,y)\land\exists^{\geq 2}x\left(E(y,x)\land\text{Red}(x)\right)\right)$
queries if $x$ is blue and it has at least one neighbor with two red neighbors. Then $\gamma$ is in GC2. Now,
$\delta(x):=\text{Blue}(x)\land\exists y\left(\lnot E(x,y)\land\exists^{\geq 2}x\left(E(y,x)\land\text{Red}(x)\right)\right)$
@@
$\eta(x):=\lnot\left(\exists y\left(E(x,y)\land\exists^{\geq 2}x\left(E(y,x)\land\text{Blue}(x)\right)\right)\right)$
@@
(Khovanskiĭ, 1991) Let $U\subseteq\mathbb{R}^{n}$ be an open set. A Pfaffian function $\sigma:U\rightarrow\mathbb{R}$ is an analytic function for which there is $\alpha\geq 1$ and a chain of analytic real functions $f_{1},\cdots,f_{r}:U\rightarrow\mathbb{R}$ such that $f_{r}=\sigma$ and that satisfies the following differential equations
$\partial f_{i}/\partial x_{j}=P_{i,j}(x,f_{1}(x),\ldots,f_{i}(x)))$
where $P_{i,j}\in\mathbb{R}[x_{1},...,x_{n},y_{1},...,y_{i}]$ are polynomials on $(n+i)$ variables of degree at most $\alpha$.